OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. 4x – 12y = 2; (10, –1)

OpenStudy (anonymous):

First find the slope of 4x - 12y = 2. What is the slope?

OpenStudy (anonymous):

the slope is 2

OpenStudy (anonymous):

Are you sure the slope is 2 for 4x - 12y = 2 ?

OpenStudy (anonymous):

A. y = 3x + 29 B. y =-3 x + 29 C. y = -1/3 + 29 D. y = -1/3 x + 7 what would be my answer

OpenStudy (anonymous):

these are my choices

OpenStudy (anonymous):

4x - 12y = 2 -12y = -4x + 2 \[ -\frac{-12y}{-12} = \frac{-4}{-12} + \frac{2}{12} \] what is the slope now?

OpenStudy (anonymous):

Do you want help or just the answer?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

Ok, what is the slope of 4x - 12y = 2?

OpenStudy (anonymous):

the slope is 3 cuz you divide 2 into 4 right

OpenStudy (anonymous):

or 2

OpenStudy (anonymous):

4x - 12y = 2 -12y = -4x + 2 \[ -\frac{-12y}{-12} = \frac{-4}{-12} + \frac{2}{12} \] Here is the equation. What is the slope?

OpenStudy (anonymous):

slope is 6

OpenStudy (anonymous):

4x - 12y = 2 -12y = -4x + 2 \[ -\frac{-12y}{-12} = \frac{-4}{-12}x + \frac{2}{12} \] equation of a line is y = mx + b m = slope. what is the slope now?

OpenStudy (anonymous):

IN the equation m = slope. \( \frac{-4}{-12} \) <---- Reduce and that is the slope of 4x - 12 = 2

OpenStudy (anonymous):

i still dont get it is the slope not 6

OpenStudy (anonymous):

can you just send me the answer please lol

OpenStudy (anonymous):

Ok, the equation we are working with is as followed 4x - 12y = 2 It is in standard form. We are going to put it into slope intercept form. We do this by using the Addition property, which we will add a -4x to eliminate it on one side and bring it over to the other side of the equation. -4x + 4x - 12y = -4x + 2 -12y = -4x + 2 Now we need to get y all by it self. We do this by using the Multiplication property. \[ \frac{-12y}{-12} = \frac{-4}{-12} + \frac{2}{12} \] \[ \frac{-12y}{-12} = \frac{-4}{-12}x + \frac{2}{12} \] \[ y = \frac{-4}{-12}x + \frac{2}{12} \] Reduce fractions \[ y = \frac{1}{3}x + \frac{1}{6} \] Do you see now and know the slope?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Once you have the slope, which is \( \frac{1}{3} \), the answer is the one with the the opposite sign and is the reciprocal of \( \frac{1}{3} \) Do you know the reciprocal of \( \frac{1}{3} \)

OpenStudy (anonymous):

honestlyt not really would that be the 3rd choice

OpenStudy (anonymous):

To find the reciprocal of \( \frac{1}{3} \) just flip is \( \frac{3}{1} \) which = 3 and then put an opposite sign and you get the slope of -3. So your answer has a slope of - 3. Do you see one of your answer with that?

OpenStudy (anonymous):

y = -3x + 29

OpenStudy (anonymous):

Yes, that is the only one with -3 as the slope.

OpenStudy (anonymous):

ohh ok that makes sense